# Moments and the Basic Equilibrium Equations

### v2

@@ Abstract.
We develop the classical theory of moments of configurations of weighted points with a focus on systems with an identically vanishing first moment. The latter condition provides equations for equilibrium configurations of systems of interacting particles under the sole condition that interactions are pairwise and along the line determined by each pair of particles. Complying external forces are admitted, so the description of some dynamical equilibrium configurations, such as relative equilibria, is included in our approach. Moments provide a unified framework for equilibrium problems in arbitrary dimensions. The equilibrium equations are homogeneous, of relatively low degree, invariant by oriented isometries (for interactions depending only on mutual distances), and are obtained through simple algebraic procedures requiring neither reduction by isometries nor a variational principle for their determination. Our equations include the renowned sets of $n$-body central configuration equations by O. Dziobek, and by A. Albouy and A. Chenciner. These equations are extended to a rather broad class of equilibrium problems, and several new sets of equilibrium equations are introduced. We also apply moments to establish a theory of constraints for mutual distances of configurations of fixed dimension and of co-spherical configurations, thus re-obtaining and adding to classical results by A. Cayley, among others.
@@ 1. Introduction
In the seminal paper [16], O. Dziobek used mutual distances as coordinates to describe central configurations as critical points of a modified potential function subject to constraints that guaranteed the configuration was planar. The constraints used by Dziobek where expressed in terms of the determinants introduced by A. Cayley [10], which later became known as Cayley-Menger determinants [8, 9]. For the planar four-body problem, Dziobek produced relatively simple algebraic equations expressing the reciprocal cubed mutual distances $r_{ij}^{-3}$ as an affine function of the product of suitably defined (volume) barycentric coordinates, $\Delta_{i}\Delta_{j}$, divided by the product of the masses, $m_{i}m_{j}$. His approach was adopted and developed by Albouy in the previously mentioned classification of four-equal-mass central configurations [1, 2]. Later on, Moeckel [29] adapted Dziobek’s method to dimension n2n-2, and, using barycentric coordinates, Albouy [4] provided a simple algebraic path to the Moeckel-Dziobek equations which did not require the (typically large) Cayley-Menger determinant.
@@
Going back to the topic of central configurations, it is worth noting that Dziobek equations also found an application in the pioneering resolution of the finiteness problem for gravitational relative equilibrium configurations [22], as well as for stationary configurations of vortices [23]. However, the main formulation of these finiteness problems actually employed a different set of central configuration equations, namely the one introduced by A. Albouy and A. Chenciner in the celebrated article [3]. Briefly, the finiteness proofs in [22, 23] applied ideas from Algebraic Geometry to exclude the possibility of a continuum of solutions, taking advantage of the fact that the so-called Albouy-Chenciner equations are algebraic. A remarkable feature of the Albouy-Chenciner equations is that they depend only on the mutual distances of the configuration, which, besides eliminating the rotational and translational symmetries, makes them at once valid for configurations of all dimensions.
@@
Unlike the known deductions of the Albouy-Chenciner equations [3, 22], which use reduction by isometries, early on in the paper we obtain these equations for every system with $\mu_{1}=\overrightarrow{O}$ as a nearly straightforward consequence of the basic interplay of the moments $\mu_{0}$ and $\mu_{1}$ with the second moment $\mu_{2}$. Later on, we provide another set of algebraic equations exhibiting the properties of the Albouy-Chenciner equations mentioned in the previous paragraph. We call these new equations the extended Leibniz identities, as they are closely related to a well-known identity involving moments and mutual distances attributed to Leibniz. We stress that, just as for the generalized Dziobek approach, the applications of the generalized Albouy-Chenciner and the extended Leibniz identities include all equilibrium problems satisfying the weaker fundamental hypothesis (WFC) stated on page 1. Moreover, concerning Geometry, the interplay of the first three moments allows us to establish a theory of constraints for distances from the points of a configuration to an arbitrary point in the ambient space, and for mutual distances between the points of a configuration of prescribed dimension, or lying on a sphere. In particular, Cayley-Menger determinantal conditions for the dimension are derived, and the number of independent Cayley-Menger constraints needed to ensure the configuration has a given dimension is expressed in terms of the configuration’s codimension $c$ simply as $c+1\choose 2$}. Our formula for the number of constraints generalizes the ones for dimensions 1, 2 and 3 in Wintner [34], $\S 357$. It underscores once again the importance of the codimension in the study of configurations of points in Euclidean affine spaces.
@@
Before moving on to a summary of specific contents, we make a brief note related to the moment problem mentioned in the first paragraph. We notice that the moment problem suggests an inverse of the equilibrium problem treated in the present work, viz., assuming the configuration $X$ is known, one would be asked to find the interactions $w_{x}$, for each $x\in X$, which make $X$ an equilibrium. For this problem, our equilibrium equations seem to offer relatively simple conditions to determine the $w_{x}$. In particular, the interactions can be determined using the linear systems provided by equations (2.2) or (4.7). If the interactions are symmetric, i.e., if Newton’s third law of motion (popularly “Action == (1)(-1) $\times$ Reaction”) holds, then, using the generalized Dziobek equations of section 6, the $w_{x}$ can be directly expressed in terms of barycentric coordinates (or mutual distances) and $c+1\choose 2$} parameters, where $c$ is the codimension of $X$.
@@
Section 6 focuses on generalizing the set of equations found by Dziobek [16] to describe central configurations as constrained critical points. Dziobek’s equations assume their simplest form when $c=1$, and, in section 6, using matrices, we extend Albouy’s arguments [4] to any symmetric family of $n$ weight vectors in $\mathbb{W}_{0}(X)$ on a common configuration $X$ of arbitrary dimension. Symmetry here means the interactions obey Newton’s third law of motion, which, in the notation of section 3, translates simply as $\varphi_{x,y}=\varphi_{y,x}$, for all $x,y\in X$.
@@ 2. Moments of a Weighted System
We begin with a brief overview of affine geometry; see [8] for a detailed exposition. Let $(\mathbb{A},\overrightarrow{\mathbb{A}},\tau)$ be an Euclidean affine space, that is, a triple formed by a point set $\mathbb{A}$, a vector space $\overrightarrow{\mathbb{A}}$ over a field $K$ endowed with an inner product $\cdot$ , and a free action $\tau$ of the additive group of $\overrightarrow{\mathbb{A}}$ on $\mathbb{A}$. The vector $\overrightarrow{v}\in\overrightarrow{\mathbb{A}}$ acts on the point $x\in\mathbb{A}$ by translating it to a point $y\in\mathbb{A}$ and is unique in that sense, so we write
@@
$G=q_{0}+\frac{1}{\mu_{0}}\sum_{x\in X}w(x)\overrightarrow{q_{0}x},$
where $q_{0}$ is an arbitrary point of $\mathbb{A}$. We call $G$ the barycenter of $(X,w)$ and denote it by $\text{bar}(X,w)$. Using the barycenter makes it possible to simplify the expressions of the first and second moments, as shown by proposition 2.3 and the Huygens-Leibniz-König theorem below.
@@
    If $\mu_{0}\neq 0$ and $G=\text{bar}(X,w)$, then $\mu_{1}(p)=\mu_{0}\overrightarrow{pG}.$
@@
    if $\mu_{0}\neq 0$ and $G=\text{bar}(X,w)$, then $\mu_{2}(p)-\mu_{2}(G)=\mu_{0}\,\overrightarrow{pG}^{2}$, for all $p\in\mathbb{A}$,
@@
If $\mu_{1}=\overrightarrow{O}$, then $\mu_{2}$ is constant according to part (2) of the HLK theorem. Thus we have (i) $\Rightarrow$ (ii); (ii) $\Rightarrow$ (iii) is immediate. If the restriction of $\mu_{2}$ to $X$ is constant and $\mu_{1}=\overrightarrow{v}$ then, from the HLK theorem,
$0=\mu_{2}(y)-\mu_{2}(x)=\overrightarrow{v}\cdot\overrightarrow{xy},\quad\text{for all}\ x,y\in X.$
@@
Corollary 2.6 has the significant implication below.
@@
We have that $\mu_{1}=\overrightarrow{O}$ if and only if $\mu_{0}=0$ and
(2.2) $\sum_{z\in X}w(z)(\overrightarrow{xy}^{2}-\overrightarrow{yz}^{2}+\overrightarrow{zx}^{2})=0,\quad\text{for every}\ x,y\in X.$
@@
Suppose $\mu_{1}=\overrightarrow{O}$, so $\mu_{0}=0$ by corollary 2.4. Thus, given $x\in X$, we can write $w(x)=-\sum_{z\in X\setminus\{x\}}w(z)$. Corollary 2.6 says that $\mu_{2}$ is constant, so for all $x\in X$, $y\in\mathbb{A}$, the equation
$\sum_{z\in X}w(z)\overrightarrow{yz}^{2}=\sum_{z\in X\setminus\{x\}}w(z)(\overrightarrow{yz}^{2}-\overrightarrow{xy}^{2})=\sum_{z\in X}w(z)(\overrightarrow{yz}^{2}-\overrightarrow{xy}^{2}),$
holds. Formulæ (2.2) follow by setting $y=x$ in the former summation and picking $y\in X$ arbitrary in the latter.
Conversely, suppose that $\mu_{0}=0$ and that formulæ (2.2) hold for all $x,y\in X$. Reversing the argument of the previous paragraph, we conclude that $\mu_{2}$ is constant on $X$. Therefore $\mu_{1}=\overrightarrow{O}$, by corollary 2.6. ∎
@@ 3. Equilibrium Equations
A configuration $X$ is an equilibrium of a systems of particles which interact according to equations (3.1) if and only if there holds either of the following sets of identities
(3.3) $\sum_{z\in X}\varphi_{x,z}(\overrightarrow{xy}^{2}-\overrightarrow{yz}^{2}+\overrightarrow{zx}^{2})=0,\quad\text{for all }x,y\in X,$
@@
(3.4) $\sum_{\{z,w\}\subset X}\varphi_{x,z}\varphi_{y,w}\overrightarrow{zw}^{2}=0,\quad\text{for all }x,y\in X.$
@@ 4. The space of weighted systems with $\mu_{1}={\overrightarrow{O}}$
The main objective is this section is to describe $\mathbb{W}_{0}(X)$.
@@
$\iota_{l}:\mathbb{W}_{0}(D^{(l)})\hookrightarrow\mathbb{W}_{0}(X),$
@@
The (oriented) volume of a configuration $X$ of codimension $c$ is the determinant $\det(\mathcal{X})$ of its configuration matrix $\mathcal{X}$ if $c=0$, and zero otherwise. Let $X_{j_{1}\ldots j_{\beta}}$ denote the subconfiguration $X\setminus\{x_{j_{1}},\ldots,x_{j_{\beta}}\}$, where $j_{1}<\ldots<j_{\beta}$, and let $\mathcal{X}_{j_{1}\ldots j_{\beta}}$ be the configuration matrix of $X_{j_{1}\ldots j_{\beta}}$. Denote the volume of $X_{j_{1}\ldots j_{\beta}}$ by $\Delta_{j_{1}\ldots j_{\beta}}$. The construction of bases for $\mathbb{W}_{0}(X)$ using the volume barycentric coordinates $\Delta_{j_{1}\ldots j_{\beta}}$ is similar to the one done in the proof of proposition 4.3. For the sake of concreteness, we provide an example.
@@
Each of the above vectors corresponds to a Dziobek subconfiguration $\{x_{1},\ldots,x_{d+1},$ $x_{l}\}$, with $l=d+2,\ldots,n$ ; we denote the weight vector associated to $l$ by $W_{l}$.
@@ 5. Some Applications to central configurations
Following [4], we say that $(X,m)$ is a central configuration if there exist a vector $\overrightarrow{\gamma}_{O}$, a point $x_{O}$ and a real number $\lambda$ such that
@@
Albouy [4] has shown that, if $a=0$, then any configuration is central, and if $a\neq 0$ and d=n1d=n-1, i.e, if $c=0$, then $X$ is a regular simplex for all values of $\mu_{0}$. In the cases $\mu_{0}\neq 0$, this assertion follows from lemma 4.1 and theorem 5.2 below (as in [4], we tacitly assume that $\lambda/\mu_{0}>0$).
@@
Multiplying both sides of (5.1) by $m_{i}$ and summing over $i$, we obtain an equation relating the constants $\lambda,x_{O},\overrightarrow{\gamma}_{O}$ and the total mass $\mu_{0}$, namely
@@
    $x_{j}=\text{bar}\left(X_{j},m\big{|}_{X_{j}}\right)$, for every $j=1,\ldots,n$.
@@
A full classification of a special family of central configurations with $\mu_{0}=0$ in the case $n=4$, including examples of central configurations with $\lambda\neq 0$, can be found in [11].
@@
A configuration of particles $(X,m)$ with $\mu_{0}\neq 0$ is central if and only if
$\sum_{i\neq j}m_{i}S_{ij}(s_{ij}+s_{jk}-s_{ki})=0,\quad\text{for every }\ j,k=1,\ldots,n.$
These formulæ are due to the authors of the celebrated reference [3] and have since been named after them. Another deduction of the Albouy-Chenciner equations can be found in [22].
@@
$(-1)^{j-d}m_{i}S_{ij}\Delta_{d+2\ldots n}=\sum_{l=d+2}^{n}m_{l}S_{lj}\Delta_{j\,d+2\ldots\widehat{l}\ldots n},\quad i=1,\ldots,d+1,\ j=1,\ldots,n.$

### v3

@@ Abstract.
We develop the classical theory of moments of configurations of weighted points with a focus on systems with an identically vanishing first moment. The latter condition produces equations for equilibrium configurations of systems of interacting particles under the sole condition that interactions are pairwise and along the line determined by each pair of particles. Complying external forces are admitted, so the description of some dynamical equilibrium configurations, such as relative equilibria in Celestial Mechanics, is included in our approach. Moments provide a unified framework for equilibrium problems in arbitrary dimensions. The equilibrium equations are homogeneous, of relatively low degree, invariant by oriented isometries (for interactions depending only on mutual distances), and are obtained through simple algebraic procedures requiring neither reduction by isometries nor a variational principle for their determination. Our equations include the renowned sets of $n$-body central configuration equations by O. Dziobek, and by A. Albouy and A. Chenciner. These equations are extended to a rather broad class of equilibrium problems, and several new sets of equilibrium equations are introduced. We also apply moments to establish a theory of constraints for mutual distances of configurations of fixed dimension and of co-spherical configurations, thus re-obtaining and adding to classical results by A. Cayley, among others.
@@ 1. Introduction
In the seminal paper [16], O. Dziobek used mutual distances as coordinates to describe central configurations as critical points of a modified potential function subject to constraints that guaranteed the configuration was planar. The constraints used by Dziobek were expressed in terms of the determinants introduced by A. Cayley [10], which later became known as Cayley-Menger determinants [8, 9]. For the planar four-body problem, Dziobek produced relatively simple algebraic equations expressing the reciprocal cubed mutual distances $r_{ij}^{-3}$ as an affine function of the product of suitably defined (volume) barycentric coordinates, $\Delta_{i}\Delta_{j}$, divided by the product of the masses, $m_{i}m_{j}$. His approach was adopted and developed by Albouy in the previously mentioned classification of four-equal-mass central configurations [1, 2]. Later on, Moeckel [29] adapted Dziobek’s method to dimension n2n-2, and, using barycentric coordinates, Albouy [4] provided a simple algebraic path to the Moeckel-Dziobek equations which did not require the (typically large) Cayley-Menger determinant.
@@
Going back to the topic of central configurations, it is worth noting that Dziobek equations also found an application in the pioneering resolution of the finiteness problem for gravitational relative equilibrium configurations [22], as well as for stationary configurations of vortices [23]. However, the main formulation of these finiteness problems employed a different set of central configuration equations, namely the one introduced by A. Albouy and A. Chenciner in the celebrated article [3]. Briefly, the finiteness proofs in [22, 23] applied ideas from Algebraic Geometry to exclude the possibility of a continuum of solutions, taking advantage of the fact that the so-called Albouy-Chenciner equations are algebraic. A remarkable feature of the Albouy-Chenciner equations is that they depend only on the mutual distances of the configuration which, besides eliminating the rotational and translational symmetries, makes them at once valid for configurations of all dimensions.
@@
Unlike the known deductions of the Albouy-Chenciner equations [3, 22], which use reduction by isometries, early on in the paper we obtain these equations for every system with $\mu_{1}=\overrightarrow{O}$ as a nearly straightforward consequence of the basic interplay of the moments $\mu_{0}$ and $\mu_{1}$ with the second moment $\mu_{2}$. Later on, we provide another set of algebraic equations exhibiting the properties of the Albouy-Chenciner equations mentioned in the previous paragraph. We call these new equations the extended Leibniz identities, as they are closely related to a well-known identity involving moments and mutual distances attributed to Leibniz. We stress that, just as for the generalized Dziobek approach, the applications of the generalized Albouy-Chenciner and the extended Leibniz identities include all equilibrium problems satisfying the weaker fundamental hypothesis (WFC) stated on page 1. Moreover, concerning Geometry, the interplay of the first three moments allows us to establish a theory of constraints for distances from the points of a configuration to an arbitrary point in the ambient space, and for mutual distances between the points of a configuration of prescribed dimension or lying on a sphere. In particular, Cayley-Menger determinantal conditions for the dimension are derived, and the number of independent Cayley-Menger constraints needed to ensure the configuration has a given dimension is expressed in terms of the configuration’s codimension $c$ simply as $c+1\choose 2$}. Our formula for the number of constraints generalizes the ones for dimensions 1, 2 and 3 in Wintner [34], $\S 357$. It underscores once again the importance of the codimension in the study of configurations of points in Euclidean affine spaces.
@@
Before moving on to a summary of specific contents, we make a brief note related to the moment problem mentioned in the first paragraph. We notice that the moment problem suggests an inverse of the equilibrium problem treated in the present work, viz., assuming the configuration $X$ is known, one would be asked to find the interactions $w_{x}$, $x\in X$, which make $X$ an equilibrium. For this problem, our equilibrium equations seem to offer relatively simple conditions to determine the $w_{x}$. In particular, the interactions can be determined using the linear systems provided by equations (2.2) or (4.3). If the interactions are symmetric, i.e., if Newton’s third law of motion (popularly “Action == (1)(-1) $\times$ Reaction”) holds, then, using the generalized Dziobek equations of section 6, the $w_{x}$ can be directly expressed in terms of barycentric coordinates (or mutual distances) and $c+1\choose 2$} parameters, where $c$ is the codimension of $X$.
@@
Section 6 focuses on generalizing the set of equations found by Dziobek [16] to describe central configurations as constrained critical points. Dziobek’s equations assume their simplest form when $c=1$, and, in section 6, using matrices, we extend Albouy’s arguments [4] to any symmetric family of $n$ weight vectors in $\mathbb{W}_{0}(X)$ on a common configuration $X$ of arbitrary dimension. For systems of interacting particles, symmetry means the interactions obey Newton’s third law of motion, which, in the notation of section 3, translates simply as $\varphi_{x,y}=\varphi_{y,x}$, for all $x,y\in X$.
@@ 2. Moments of a Weighted System
We begin with a brief overview of affine geometry; see [8] for a thorough exposition. Let $(\mathbb{A},\overrightarrow{\mathbb{A}},\tau)$ be an Euclidean affine space, that is, a triple formed by a point set $\mathbb{A}$, a vector space $\overrightarrow{\mathbb{A}}$ over a field $K$ endowed with an inner product $\cdot$ , and a free action $\tau$ of the additive group of $\overrightarrow{\mathbb{A}}$ on $\mathbb{A}$. The vector $\overrightarrow{v}\in\overrightarrow{\mathbb{A}}$ acts on the point $x\in\mathbb{A}$ by translating it to a point $y\in\mathbb{A}$ and is unique in that sense, so we write
@@
$G=q+\frac{1}{\mu_{0}}\sum_{x\in X}w(x)\overrightarrow{qx},$
where $q$ is an arbitrary point of $\mathbb{A}$. We call $G$ the barycenter of $(X,w)$ and denote it by $\text{bar}(X,w)$. Using the barycenter makes it possible to simplify the expressions of the first and second moments, as shown by proposition 2.3 and the Huygens-Leibniz-König theorem below.
@@
    If $\mu_{0}\neq 0$ and $G=\text{\emph{bar}}(X,w)$, then $\mu_{1}(p)=\mu_{0}\overrightarrow{pG}.$
@@
    if $\mu_{0}\neq 0$ and $G=\text{\emph{bar}}(X,w)$, then $\mu_{2}(p)-\mu_{2}(G)=\mu_{0}\,\overrightarrow{pG}^{2}$, for all $p\in\mathbb{A}$,
@@
If $\mu_{1}=\overrightarrow{O}$, then $\mu_{2}$ is constant according to part (2) of the HLK theorem. Thus we have (i) $\Rightarrow$ (ii); (ii) $\Rightarrow$ (iii) is immediate. If the restriction of $\mu_{2}$ to $X$ is constant and $\mu_{1}=\overrightarrow{v}$ then, from the HLK theorem, part (2),
$0=\mu_{2}(y)-\mu_{2}(x)=2\overrightarrow{v}\cdot\overrightarrow{xy},\quad\text{for all}\ x,y\in X.$
@@
Corollary 2.6 has the significant implications below.
@@
We have that $\mu_{1}=\overrightarrow{O}$ if and only if $\mu_{0}=0$ and either of the equivalent equations
(2.2) $\sum_{z\in X}w(z)(\overrightarrow{xz}^{2}-\overrightarrow{yz}^{2})=0,\quad\text{or}\quad\sum_{z\in X\setminus\{x\}}w(z)(\overrightarrow{xy}^{2}-\overrightarrow{yz}^{2}+\overrightarrow{zx}^{2})=0$
holds for every $x,y\in X$.
@@
Equations (2.2) are mere restatements of assertion (iii) in corollary 2.6. The equations are equivalent since
$\sum_{z\in X}w(z)(\overrightarrow{xz}^{2}-\overrightarrow{yz}^{2})=\sum_{z\in X\setminus\{x\}}w(z)(\overrightarrow{xz}^{2}-\overrightarrow{yz}^{2})+\left(\sum_{z\in X\setminus\{x\}}w(z)\right)\overrightarrow{xy}^{2}.$
@@ 3. Equilibrium Equations
A configuration $X$ is an equilibrium of a systems of particles which interact according to equations (3.1) if and only if, for every pair of points $x,y\in X$, there holds one of the following identities
(3.3) $\sum_{z\in X\setminus\{x\}}\varphi_{x,z}(\overrightarrow{xy}^{2}-\overrightarrow{yz}^{2}+\overrightarrow{zx}^{2})=0,$
@@
(3.4) $\sum_{\{z,w\}\subset X}\varphi_{x,z}\varphi_{y,w}\overrightarrow{zw}^{2}=0.$
@@ 4. The space of weighted systems with $\mu_{1}={\overrightarrow{O}}$
The main objective in this section is to describe $\mathbb{W}_{0}(X)$.
@@
(4.1) $\iota_{l}:\mathbb{W}_{0}(D^{(l)})\hookrightarrow\mathbb{W}_{0}(X),$
@@
The (oriented) volume of a configuration $X$ of codimension $c$ is the determinant $\det(\mathcal{X})$ of its configuration matrix $\mathcal{X}$ if $c=0$, and zero otherwise. Let $X_{j_{1}\ldots j_{\beta}}$ denote the subconfiguration $X\setminus\{x_{j_{1}},\ldots,x_{j_{\beta}}\}$, where $j_{1}<\ldots<j_{\beta}$, and let $\mathcal{X}_{j_{1}\ldots j_{\beta}}$ be the configuration matrix of $X_{j_{1}\ldots j_{\beta}}$. Denote the volume of $X_{j_{1}\ldots j_{\beta}}$ by $\Delta_{j_{1}\ldots j_{\beta}}$. The construction of bases for $\mathbb{W}_{0}(X)$ using the volume barycentric coordinates $\Delta_{j_{1}\ldots j_{\beta}}$ is similar to the one done in the proof of proposition 4.3 and employs the embeddings $\iota_{l}$, see (4.1). For the sake of concreteness, we provide an example.
@@
Each of the above weight vectors corresponds to a Dziobek subconfiguration $\{x_{1},\ldots,x_{d+1},$ $x_{l}\}$, with $l=d+2,\ldots,n$ ; we denote the weight vector associated to $l$ by $W_{l}$.
@@ 5. Some Applications to central configurations
Following Albouy [4], we say that $(X,m)$ is a central configuration if there exist a vector $\overrightarrow{\gamma}_{O}$, a point $x_{O}$ and a real number $\lambda$ such that
@@
Albouy has shown that, if $a=0$, then any configuration is central, and if $a\neq 0$ and d=n1d=n-1, i.e, if $c=0$, then $X$ is a regular simplex for all values of $\mu_{0}$. In the cases $\mu_{0}\neq 0$, this assertion follows from lemma 4.1 and theorem 5.2 below (as in [4], we tacitly assume that $\lambda/\mu_{0}>0$).
@@
Multiplying both sides of (5.1) by $m_{j}$ and summing over $j$, we obtain an equation relating the constants $\lambda,x_{O},\overrightarrow{\gamma}_{O}$ and the total mass $\mu_{0}$, namely
@@
    $x_{j}=\text{\emph{bar}}\left(X_{j},m\big|_{X_{j}}\right)$, for every $j=1,\ldots,n$.
@@
In the case $n=4$, a full classification of a special family of central configurations with $\mu_{0}=0$, including examples of central configurations with $\lambda\neq 0$, can be found in [11].
@@
A configuration of particles $(X,m)$ with $\mu_{0}\neq 0$ is central if and only if, for each $j,k=1,\ldots,n$, there holds
$\sum_{i\neq j}m_{i}S_{ij}(s_{ij}+s_{jk}-s_{ki})=0.$
The above formulæ  were found by the authors of the celebrated reference [3] and have since been named after them. Reference [22] contains an alternative deduction.
@@
$(-1)^{i-d}m_{i}S_{ij}\Delta_{d+2\ldots n}=\sum_{l=d+2}^{n}m_{l}S_{lj}\Delta_{i\,d+2\ldots\widehat{l}\ldots n},\quad i=1,\ldots,d+1,\ j=1,\ldots,n.$